# Adaptive Group Robust Ensemble Knowledge Distillation

### v1

@@ Abstract
Neural networks can learn spurious correlations in the data, often leading to performance disparity for underrepresented subgroups. Studies have demonstrated that the disparity is amplified when knowledge is distilled from a complex teacher model to a relatively “simple” student model. Prior work has shown that ensemble deep learning methods can improve the performance of the worst-case subgroups; however, it is unclear if this advantage carries over when distilling knowledge from an ensemble of teachers, especially when the teacher models are debiased. This study demonstrates that traditional ensemble knowledge distillation can significantly drop the performance of the worst-case subgroups in the distilled student model even when the teacher models are debiased. To overcome this, we propose Adaptive Group Robust Ensemble Knowledge Distillation (AGRE-KD), a simple ensembling strategy to ensure that the student model receives knowledge beneficial for unknown underrepresented subgroups. Leveraging an additional biased model, our method selectively chooses teachers whose knowledge would better improve the worst-performing subgroups by upweighting the teachers with gradient directions deviating from the biased model. Our experiments on several datasets demonstrate the superiority of the proposed ensemble distillation technique and show that it can even outperform classic model ensembles based on majority voting.
@@ 1 Introduction
When trained with empirical risk minimization (ERM), neural networks are susceptible to capturing spurious correlations in the data (Tiwari and Shenoy,, 2023), which are features that correlate with but not causally related to the class label (Qiu et al.,, 2023). In particular, the class label might spuriously correlate with patterns in the data that are easier to learn than the intended pattern. For example, in the Waterbirds dataset (Sagawa et al.,, 2019), which contains images of landbirds and waterbirds, most landbirds images have a land background, and waterbirds images have a water background. Instead of predicting the actual bird species, the model trained with ERM can achieve high accuracy by looking at the background. This results in a significantly higher error for underrepresented subgroups that do not exhibit the spurious correlation (e.g., landbird on water background and water bird on land background). Several works have shown that model compression methods such as pruning (Hooker et al.,, 2020), and knowledge distillation (Lukasik et al.,, 2023; Lee and Lee,, 2023; Wang et al.,, 2023) can exacerbate the performance disparities between different subgroups. In knowledge distillation (KD), a network with a smaller capacity (student) is trained using the output of a higher capacity network (teacher) (Hinton et al.,, 2015). While KD can improve the student model’s average performance, the gain is not uniform across subgroups (Lukasik et al.,, 2023).
@@
On the other hand, deep ensemble models have been shown to enhance generalization performance compared to individual models (Ganaie et al.,, 2022), and simple deep ensemble models with the same architecture, objective, and optimization settings can attenuate this shortcoming and improve the worst-case group performance (Ko et al.,, 2023; Kenfack et al.,, 2021). However, evaluating several models at test time can be computationally expensive, making them less practical for deployment on edge devices. To address this issue, ensemble knowledge distillation involves distilling the knowledge of multiple teachers to a single student model (You et al.,, 2017; Radwan et al.,, 2024), and it remains unclear whether distilling from an ensemble of teachers improves student’s worst-group performance.
@@
This paper studies how knowledge distillation from multiple teachers impacts underrepresented subgroups. We investigate whether the subgroup performance gains of deep ensemble models apply to ensemble knowledge distillation. Focusing on logit distillation, we consider teacher models debiased by the last-layer retraining  (Kirichenko et al.,, 2022) and investigate whether the student model can learn debiased representations from the output of retrained last-layer of the teacher. In last-layer retraining, a small held-out validation set of the group-balanced data is used to retrain the last layer of the teacher model to mitigate the spurious correlation. Our results reveal that underrepresented subgroups can be negatively impacted when distilling from multiple teachers, even when the teachers are debiased. There are other ensemble distillation approaches designed to boost the student’s performance by modeling a better aggregation of the teachers’ knowledge (Du et al.,, 2020; Zhang et al.,, 2022). In particular, the ensemble distillation method proposed by  Du et al., (2020) aims to find a better compromise when teachers have conflicting predictions, and the method by  Zhang et al., (2022) ensures distillation is done only using teachers with confident predictions. Our results show that these ensemble distillation methods cannot effectively fix the performance disparity of the student.
@@
We propose an Adaptive Group Robust Ensemble Knowledge Distillation (AGRE-KD) method to encourage the student to improve the performance of unknown worst-case subgroups. Specifically, our method relies on a model that has captured the spurious correlation (i.e., a biased model) to guide the teachers’ outputs aggregation process and ensure the student model does not capture biased knowledge from the teacher. Prior work has relied on a reference classifier to target and upweight samples from unknown worst-case, using the errors of the reference classifier (Liu et al.,, 2021; Nam et al.,, 2020) or its per-sample gradient magnitude (Ahn et al.,, 2022). In contrast, our proposal uses the gradient direction of the biased model to select and weigh teachers’ outputs adaptively during the training.
@@
Intuitively, training the student model solely following the gradient direction that minimizes its KD loss with a biased teacher can result in local/global minima with a higher loss for the underrepresented. Our results suggest this behavior can be exacerbated in ensemble knowledge distillation since the inherent consensus from the teachers’ gradient direction can be dominated by a direction that minimizes the average error at the expense of the worst-case error. Our proposed methods compare the gradient direction of student loss with the biased model and each teacher in the ensemble and upweight teachers whose gradient direction deviates the most from the biased model since minimizing the student loss with the biased model likely results in poor worst-case group error. Figure 1 illustrates the gradient directions of the student loss with the biased model and three teachers and shows how our weighting scheme ensures the aggregated gradient direction is mainly influenced by the least biased model. As the gradient magnitude can be very noisy, we compare gradient directions by computing the dot product of their normalized vectors (i.e., the cosine similarity), which removes the influence of the gradient magnitude. The contribution of this paper can be summarized as follows:
@@
  • We demonstrate empirically that ensemble knowledge distillation amplifies performance disparities contrary to deep ensemble models. We attribute this to the reduced capacity of the student network by showing that ensemble self-distillation using models with the same capacity reduces the performance disparity but, in some cases, achieves comparable performance with the teachers.
@@
  • We propose a novel gradient-based weighting scheme to ensure the student model minimizes teachers’ loss towards better worst-case group error. The proposed method acts in the gradient space and utilizes a model that has learned the spurious correlation (biased model) to orchestrate the distillation process.
@@
  • We perform intensive experiments on three well-known benchmarks, and the results demonstrate the superiority of the proposed method.
@@ 2 Background
We consider a multiclass classification task using a given the training $D=\{(x_{i},y_{i})\}_{i=1}^{m}$, where $x_{i}\in\mathcal{X}$ is input feature and $y_{i}\in\mathcal{Y}$ the target variable with $c=|\mathcal{Y}|$ classes, we aim to build a classifier $h(\cdot)$ that accurately predicts the class of the unlabelled test dataset. The classifier uses mapping function $f:\mathcal{X}\rightarrow\mathbb{R}^{c}$, that assigns scores $[\sigma_{1}(z_{1}),\dots,\sigma_{c}(z_{c})]$, such that $z$ is the output logits of the a given sample $x$ and $\sigma_{y}(z)$ the softmax function defined as $\sigma_{y}(z)=\frac{\exp(z_{y})}{\sum_{j\in[c]}\exp(z_{j})},\;\forall y\in[c]$.
The classifier is derived by predicting the class label that maximizes the softmax, $h(x)=\operatorname{argmax}_{j\in[c]}\sigma_{j}(z)$. We evaluate the classifier’s performance during training using a loss function, such as the softmax cross-entropy loss function, which measures how accurately samples are classified.
@@
In KD, a student network $f^{s}$ aims to achieve performance close to the higher-capacity network by mimicking the teacher model $f^{t}$ Hinton et al., (2015). In practice, we train the student model to mimic the teacher’s output by minimizing the Kullback-Leibler (KL) divergence between their outputs, defined as follows:
@@
where $z^{s}$ and $z^{t}$ are the student and the teacher logits, respectively. $\tau$ is the temperature parameter controlling the smoothness of the probability distribution for more fine-grained information. The student loss is combined with the classification loss on the ground truth label. The overall student loss is defined by the equation 2,
@@
where $\mathcal{L}_{\text{cls}}$ is the classification loss (e.g., cross-entropy loss) between the student’s output and the ground truth label ($y$), and $\alpha$ is a hyperparameter controlling the classification loss and knowledge distillation loss. In other KD techniques, the student network is enforced to mimic the teacher’s internal representation instead of only outputs (Romero et al.,, 2014). Transferring knowledge from intermediary representation (feature-level) can provide more fine-grained information and boost the students’ performance Romero et al., (2014). Recent studies have shown that models trained with ERM still learn core features, and spurious features are only amplified in the network’s last layer (Kirichenko et al.,, 2022; Qiu et al.,, 2023; LaBonte et al.,, 2024). In this regard, we restrict ourselves to logit distillation and leave feature distillation for future exploration.
Like in ensemble learning, distilling knowledge from multiple teachers instead of a single one is expected to improve the student model (You et al.,, 2017). Ensemble learning is a widely used technique to boost the generalization performance of a model (Allen-Zhu and Li,, 2020). Studies have shown that training several independent models and averaging their predictions at test time can yield a model that outperforms every individual model in the ensemble (Ganaie et al.,, 2022). However, evaluating multiple models for predictions at test time limits the practical use of deep ensembles due to computational overhead. Knowledge distillation can address this issue by enforcing the student network to mimic the ensemble’s output. In ensemble KD, we train the student model using the averaged softened output or the averaged knowledge distillation loss 111(Du et al.,, 2020) showed that the averaged softened output is equal to averaged KD loss. as follows:
@@
Where $M$ is the number of teachers and the loss 3 can be plugged in equation 2 for training the student with KD from an ensemble of $M$ teachers. In contrast to these methods, our method is fully unsupervised, both based in terms of ground truth class label and group information. Related works more aligned with our approach are unsupervised methods that leverage teachers’ diverse knowledge to improve students’ performance. (You et al.,, 2017; Du et al.,, 2020; Fukuda et al.,, 2017; Zhang et al.,, 2022; Kwon et al.,, 2020). For example, Fukuda et al., (2017) suggest that randomly selecting a teacher during mini-batch training can allow the student model to capture complementary knowledge of teachers. Other authors argue that simply averaging teachers’ softmax outputs can mislead the student model, particularly when there is competition or contradictions between teachers. In this regard, Zhang et al., (2022) proposes a sample-wise weighting for teacher loss based on the confidence of the teacher’s prediction compared to the ground label. Other methods consider label-free weighing schemes by comparing teachers in the gradient space. For instance, Du et al., (2020) use a multi-objective optimization approach in the gradient space to find teachers that agree the most in the gradient direction that minimizes their loss with the student model. Similarly, Zhou et al., (2024) sample-wise teachers selection by only averaging the majority of teachers with the same gradient directions. Our method also operates within gradient space to enhance students’ resilience to spurious correlation. Our experiments demonstrate that adhering to the majority of teachers’ opinions does not always benefit the underrepresented subgroups.
@@ 3 Related Work
We consider settings where samples in the dataset $D$ are associated with unknown group labels that spuriously correlate with the class label. For example, in the CelebA dataset, the class ‘hair color‘ (blond, non-blond) correlates with the gender (male, female) since most images with blond hair belong to the female group. Neural networks can capture this correlation, resulting in poor performance for certain subgroups (e.g., blond males) (Sagawa et al.,, 2019). We aim to ensure that the model does not capture spurious correlations in the data and accurately classifies samples from all subgroups. In particular, we measure the model’s bias using the performance of worst-case group. Several methods have been proposed to mitigate these biases in single models (Kenfack et al., 2024b, ). When the group information is known, Sagawa et al., (2019) propose Group Distributionally Robust Optimization (DRO) that minimizes loss of the group experiencing the maximum loss. However, group information can be costly to collect or unavailable due to privacy restrictions (Lahoti et al.,, 2020; Kenfack et al., 2024a, ). Several methods have been proposed in this setting to improve the worst-case group performance without group labels (Kenfack et al., 2024b, ). For instance, Lahoti et al., (2020) proposes to use an adversary to up-weight regions where the model makes the most mistakes and demonstrates that this adversary can upweight sample from the worst-case group. Liu et al., (2021) and Nam et al., (2020) rely on a reference classifier to a reference classifier to target and upweight worst-performing. These methods use the mistakes of the reference classifier to improve group robustness by up-weighting misclassified samples (Liu et al.,, 2021). The reference classifier is generally trained to amplify the misclassification of samples from the unknown worst-performing group (Nam et al.,, 2020). While these methods do not use group labels during the training, they require access to a small validation set with group labels for model selection or hyperparameter tuning (Kenfack et al., 2024b, ).  Kirichenko et al., (2022) proposed Deep Feature Reweighting (DRF) for training group robust model by training the model empirical risk minimization (ERM) on the training dataset, and then retraining the last layer of neural network with a small subset of a held-out group-balanced dataset. The proposed method achieves state-of-the-art worst-case group performance. The study demonstrated that neural networks can encode relevant and spurious features during training. Still, the impact of spurious features is upweighted in the classification layer due to high group imbalance. In a subsequent study, LaBonte et al., (2024); Qiu et al., (2023) demonstrated comparable performance by fine-tuning the last layer with fewer group information or without group labels using proxy group information from a reference classifier. This work builds on Kirichenko et al., (2022); LaBonte et al., (2024) for debiasing the teacher models by retraining their last layers using a held-out group balanced with fewer group annotations. We investigate whether distilling knowledge from an ensemble of (debiased) teachers can lead to more robust student models.
@@
Several works have studied bias in knowledge distillation with a single teacher model (Lukasik et al.,, 2023; Lee and Lee,, 2023; Lukasik et al.,, 2023; Tiwari et al.,, 2024; Bassi et al.,, 2024). In particular, Lukasik et al., (2023) demonstrates that teacher errors can be amplified by the student during distillation and proposed a mitigation strategy that distills only the confident predictions of the teacher. However, their study focuses on worst-class errors and KD with a single teacher, while we study worst-subgroup errors with multiple teachers. Lee and Lee, (2023) propose an adapted version of Simple knowledge distillation (SimKD (Chen et al.,, 2022)) that transplants the last layer of the teacher to student and only distills features. With the teacher trained with Group DRO, they show that transplanting the teacher’s last layer to the student only improves the worst-case group performance if the feature distillation is performed by upweighting misclassified sample from a reference classifier (Lee and Lee,, 2023). However, their method uses the reference classifier by Liu et al., (2021), and the efficiency of sample weighting requires intensive tuning of the number of epochs to train the reference classifier. Similarly,  Tiwari et al., (2024) uses the earlier layers of the neural network to train a reference classifier and shows that it improves the recall of worst group samples within the misclassification set, which are upweighted in the KD loss. However, their method also requires class labels to derive the misclassified samples.. Lukasik et al., (2023) study where it is best to apply the debiasing mechanism (Group DRO) and conclude that applying the robust loss to both the teacher and student model improves the average performance along with the worst-case group performance. In contrast to these prior works, we study bias in knowledge distillation with multiple teachers without group information and class labels. We investigate whether the subgroup’s performance gain observed in deep ensemble models also applies when the knowledge of the ensemble is distilled to a single model. We aim to achieve better worst-case performance across subgroups when aggregating the outputs of multiple teachers in knowledge distillation. To the best of our knowledge, this represents the first study on bias in ensemble knowledge distillation.

### v2

@@ Abstract
Neural networks can learn spurious correlations in the data, often leading to performance degradation for underrepresented subgroups. Studies have demonstrated that the disparity is amplified when knowledge is distilled from a complex teacher model to a relatively “simple” student model. Prior work has shown that ensemble deep learning methods can improve the performance of the worst-case subgroups; however, it is unclear if this advantage carries over when distilling knowledge from an ensemble of teachers, especially when the teacher models are debiased. This study demonstrates that traditional ensemble knowledge distillation can significantly drop the performance of the worst-case subgroups in the distilled student model even when the teacher models are debiased. To overcome this, we propose Adaptive Group Robust Ensemble Knowledge Distillation (AGRE-KD), a simple ensembling strategy to ensure that the student model receives knowledge beneficial for unknown underrepresented subgroups. Leveraging an additional biased model, our method selectively chooses teachers whose knowledge would better improve the worst-performing subgroups by upweighting the teachers with gradient directions deviating from the biased model. Our experiments on several datasets demonstrate the superiority of the proposed ensemble distillation technique and show that it can even outperform classic model ensembles based on majority voting. Our source code is available at https://github.com/patrikken/AGRE-KD
@@ 1 Introduction
When trained with empirical risk minimization (ERM), neural networks are susceptible to capturing spurious correlations in the data (Tiwari & Shenoy, 2023), which are features that correlate with but not are causally related to the class label (Qiu et al., 2023). In particular, the class label might spuriously correlate with patterns in the data that are easier to learn than the intended pattern. For example, in the Waterbirds dataset (Sagawa et al., 2019), which contains images of landbirds and waterbirds, most landbirds are located in a land background, and waterbirds in a water background. Instead of predicting the actual bird species, the model trained with ERM can achieve high accuracy by relying on the background to make predictions. This results in a significantly higher error for underrepresented subgroups that do not exhibit the spurious correlation (e.g., landbird on water background and waterbird on land background). Several works have shown that model compression methods such as pruning (Hooker et al., 2020), and knowledge distillation (Lukasik et al., 2023; Lee & Lee, 2023; Wang et al., 2023) can exacerbate the performance disparities between different subgroups. In knowledge distillation (KD), a network with a smaller capacity (student) is trained using the output of a higher capacity network (teacher) (Hinton et al., 2015). While KD can improve the student model’s average performance, the gain is not uniform across subgroups (Lukasik et al., 2023).
@@
On the other hand, deep ensemble models have been shown to enhance generalization performance compared to individual models (Ganaie et al., 2022), and simple deep ensemble models with the same architecture, objective, and optimization settings can attenuate this shortcoming and improve the worst-case group performance (Ko et al., 2023; Kenfack et al., 2021). However, evaluating several models at test time can be computationally expensive, making them less practical for deployment on edge devices. To address this issue, ensemble knowledge distillation involves distilling the knowledge of multiple teachers to a single student model (You et al., 2017; Radwan et al., 2024), and it remains unclear whether distilling from an ensemble of teachers improves the student’s worst-group performance.
@@
This paper studies how knowledge distillation from multiple teachers impacts underrepresented subgroups. We investigate whether the subgroup performance gains of deep ensemble models apply to ensemble knowledge distillation. Focusing on logit distillation, we consider ERM-trained (biased) and debiased teachers obtained by deep feature reweighting (DFR). DFR is a method proposed by Kirichenko et al. (2022), showing that simply retraining the last layer of the ERM-trained model using a small held-out validation set of group-balanced data can mitigate the spurious correlation. In this work, we also investigate whether the student model can learn debiased representations from the teacher’s output, whose only the last layer was retrained to mitigate bias.
The rationale behind focusing on logit distillation for investigating bias in ensemble knowledge is also motivated by the study by Izmailov et al. (2022), where the authors demonstrated that several debiasing methods do not learn better feature representation but correctly weight core features in the last classification layer, which reduce the reliance on spurious features for predictions. Our results reveal that underrepresented subgroups can be negatively impacted when training the student model using the aggregated outputs of multiple teachers. Other ensemble distillation approaches have been designed to boost the student’s performance by modelling a better aggregation of the teachers’ knowledge (Du et al., 2020; Zhang et al., 2022). In particular, the ensemble distillation method proposed by  Du et al. (2020) aims to find a better compromise when teachers have conflicting predictions, and the method by  Zhang et al. (2022) ensures distillation is done only using teachers with confident predictions. Our results show that these ensemble distillation methods do not effectively fix the performance disparity across groups in the student model.
@@
We therefore propose an Adaptive Group Robust Ensemble Knowledge Distillation (AGRE-KD) method to encourage the student to improve the performance of unknown worst-case subgroups. Specifically, our method relies on a model that has captured the spurious correlation (i.e., a biased model) to guide the teachers’ outputs aggregation process and ensure the student model does not capture biased knowledge from teachers. Prior work has relied on a reference classifier to target and upweight samples from unknown worst-case, using the errors of the reference classifier (Liu et al., 2021; Nam et al., 2020) or its per-sample gradient magnitude (Ahn et al., 2022). In contrast, our proposal uses the gradient direction of the biased model to select and weigh teachers’ outputs adaptively during the training.
@@
Intuitively, training the student model solely following the gradient direction that minimizes its KD loss with a biased teacher can result in local/global minima with a higher loss for the underrepresented groups (Lukasik et al., 2021). Our results suggest this behaviour can be exacerbated in ensemble knowledge distillation since the inherent consensus from the teachers’ gradient direction can be dominated by a direction that minimizes the average error at the expense of the worst-group error. Our proposed methods compare the gradient direction of student loss with the biased model and each teacher in the ensemble, and upweight teachers whose gradient direction deviates the most from the biased model. Figure 1 illustrates the gradient directions of the student loss with the biased model and three teachers and shows how classic averaging yields a direction that mostly aligns with the biased teachers, while our adaptive weight prioritizes the least biased models. As the gradient magnitude can be very noisy, we compare gradient directions by computing the dot product of their normalized vectors (i.e., the cosine similarity), which removes the influence of the gradient magnitude. The contributions of this paper can be summarized as follows:
@@
  • We demonstrate empirically that ensemble knowledge distillation amplifies the performance drop for the worst-case group, contrary to deep ensemble models. We attribute this to the reduced capacity of the student network by showing that ensemble self-distillation using models with the same capacity improves robustness to spurious correlation and achieves comparable performance with the teachers.
  • We show that distilling knowledge from teacher models with only their last layer retrained to mitigate bias provides significantly more robust knowledge to the student model.
@@
  • We propose a novel gradient-based weighting scheme to ensure the distillation process minimizes the student’s loss towards better worst-case group error. The proposed method utilizes a model that has learned the spurious correlation (biased model) to orchestrate the distillation process in the gradient space.
@@
  • We perform intensive experiments on four well-known benchmarks, and the results demonstrate the superiority of the proposed method.
@@ 2 Background
We consider a multiclass classification task using a given the training $D=\{(x_{i},y_{i},a_{i})\}_{i=1}^{m}$, where $x_{i}\in\mathcal{X}$ is the input feature, $y_{i}\in\mathcal{Y}$ the target variable with $c=|\mathcal{Y}|$ classes, and $a_{i}\in\mathcal{A}$ an unknown spurious feature. We aim to build a classifier $h(\cdot)$ that accurately predicts the class of the unlabelled test dataset. The classifier uses mapping function $f:\mathcal{X}\rightarrow\mathbb{R}^{c}$, that assigns scores $\sigma_{1}(z_{1}),\dots,\sigma_{c}(z_{c})]$, such that $z$ is the output logits of the given sample $x$ and $\sigma_{y}(z)$ the softmax function defined as $\sigma_{y}(z)=\frac{\exp(z_{y})}{\sum_{j\in[c]}\exp(z_{j})},\;\forall y\in[c]$.
The classifier is derived by predicting the class label that maximizes the softmax, $h(x)=\operatorname{argmax}_{j\in[c]}\sigma_{j}(z)$. We evaluate the classifier’s performance during training using a loss function, such as the softmax cross-entropy loss function, which measures how accurately samples are classified. When the attribute $a$ spuriously correlates with the target $y$, ERM-trained models can strongly rely on the spurious feature and fail to generalize on the subgroups where the spurious correlation does not hold (Ye et al., 2024).
@@
In KD, a student network $f^{s}$ aims to achieve performance close to the higher-capacity network by mimicking the teacher model $f^{t}$ (Hinton et al., 2015). In practice, we train the student model to mimic the teacher’s output by minimizing the Kullback-Leibler (KL) divergence between their outputs, defined as follows:
@@
where $z^{s}$ and $z^{t}$ are the student and the teacher logits, respectively. $\tau$ is the temperature parameter controlling the smoothness of the probability distribution for more fine-grained information. The student loss can be combined with the classification loss on the ground truth label using the following double loss,
@@
where $\mathcal{L}_{\text{cls}}$ is the classification loss (e.g., cross-entropy loss) between the student’s output and the ground truth label ($y$), and $\alpha$ is a hyperparameter controlling the classification loss and knowledge distillation loss. In other KD techniques, the student network is enforced to mimic the teacher’s internal representation instead of only outputs (Romero et al., 2014). Transferring knowledge from intermediary representation (feature-level) can provide more fine-grained information and boost the students’ performance (Romero et al., 2014). On the other hand, recent studies have shown that models trained with ERM still learn core features while spurious features are only amplified in the last classifier layer (Kirichenko et al., 2022; Qiu et al., 2023; Izmailov et al., 2022). In this regard, we restrict ourselves to logit distillation and leave feature distillation for future exploration.
Deep Ensemble.
Ensemble learning is a widely used technique to boost the generalization performance of a model (Allen-Zhu & Li, 2020). Studies have shown that training several independent models and averaging their predictions at test time can yield a model that outperforms each individual model in the ensemble (Ganaie et al., 2022). However, evaluating multiple models for predictions at test time limits the practical use of deep ensembles due to computational overhead. Knowledge distillation can address this issue by enforcing the student network to mimic the ensemble’s output.
Ensemble KD.
Like in ensemble learning, distilling knowledge from multiple teachers instead of a single one is expected to improve the student model (You et al., 2017). In ensemble KD, we train the student model using the averaged softened output or the averaged knowledge distillation losses of $M$ teachers [^1] Du et al. (2020) showed that the averaged softened output is equal to averaged KL loss. as follows:
@@
This ensemble KD loss can be plugged in equation 2 for training the student with KD from an ensemble of $M$ teachers. In contrast to these methods, our method is fully unsupervised, both based in terms of ground truth class label and group information. Related works more aligned with our approach are unsupervised methods that leverage teachers’ diverse knowledge to improve students’ performance. (You et al., 2017; Du et al., 2020; Fukuda et al., 2017; Zhang et al., 2022; Kwon et al., 2020). For example, Fukuda et al. (2017) suggest that randomly selecting a teacher during mini-batch training can allow the student model to capture complementary knowledge of teachers. Other authors argue that simply averaging teachers’ softmax outputs can mislead the student model, particularly when there is competition or contradictions between teachers. In this regard, Zhang et al. (2022) proposes a sample-wise weighting for teacher loss based on the confidence of the teacher’s prediction compared to the ground label. Other methods consider label-free weighing schemes by comparing teachers in the gradient space. For instance, Du et al. (2020) use a multi-objective optimization approach in the gradient space to find teachers that agree the most in the gradient direction that minimizes their loss with the student model. Similarly, Zhou et al. (2024) sample-wise teachers selection by only averaging the majority of teachers with the same gradient directions. Our method also operates within gradient space to enhance students’ resilience to spurious correlation. Our experiments demonstrate that adhering to the majority of teachers’ opinions does not always benefit the underrepresented subgroups.
Spurious Correlation
Let $\mathcal{D}_{tr}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{n}$ be the training dataset, where each input $x_{i}\in\mathcal{X}$ belongs to one of $K$ classes $y_{i}\in\mathcal{Y}$. Along with the label $y_{i}$, each sample also contains a spurious attribute $a_{i}\in\mathcal{A}$, which is correlated with but not causally predictive of $y_{i}$. A spurious correlation, denoted $\langle y,a\rangle$, refers to the association between a class $y\in\mathcal{Y}$ and a spurious attribute $a\in\mathcal{A}$. In practice, an attribute $a$ may be linked to multiple classes, i.e., $\phi:\mathcal{A}\mapsto\mathcal{Y}^{K^{\prime}}$ for some $1<K^{\prime}\leq K$, depending on the training data $\mathcal{D}_{tr}$. To capture this structure, each sample can be assigned a group label $g=(y,a)$, and the set of all such group labels is $\mathcal{G}=\mathcal{Y}\times\mathcal{A}$. For example, within $\mathcal{D}_{tr}$, the same attribute $a$ might appear in two different classes, giving rise to correlations $\langle y,a\rangle$ and $\langle y^{\prime},a\rangle$. A model trained naively on this data may incorrectly rely on $a$ to predict $y$, which will lead to systematic errors when encountering samples with $\langle y^{\prime},a\rangle$ (Ye et al., 2024). As the average accuracy does not fully capture the robustness of the model to spurious correlation, we use the worst-group accuracy (WGA) to measure the model’s reliance on spurious correlation for predictions. Furthermore, in the presence of spurious features, a model $f$ parametrized by $\theta$ is optimized to minimize the loss of the worst-performing subgroup, i.e.,
$\arg\min_{\theta}\max_{g\in\mathcal{G}}\frac{1}{|g|}\sum_{i\in g}^{|g|}\mathcal{L}(f(x_{i};\theta),y_{i})$ (4)
The WGA therefore quantifies the accuracy of the group experiencing the smallest test accuracy. Another line of related work aims to make the model robust to distribution shift (Yang et al., 2024), especially covariate shift. Unlike spurious correlation, covariate shift occurs when the training and test data follow different distributions, while the conditional distribution of output values given input points is unchanged. In other words, the marginal distribution of features changes, but the conditional relationship between labels and features remains unchanged. Techniques like importance weighting or domain adaptation aim to correct for this shift (Sugiyama et al., 2007), while we focus in this work on addressing model reliance on spurious features during training. Both notions could be related if the ’spurious feature’ is the source of the covariate shifts, i.e., a feature that was predictive for the training data becomes spurious in the test set; however, such scenarios are less likely to occur since the model should not rely on spurious features.
@@ 3 Related Work
We consider settings where samples in the dataset $D$ are associated with unknown group labels that spuriously correlate with the class label. For example, in the CelebA dataset, the class ‘hair color‘ (blond, non-blond) correlates with the gender (male, female) since most images with blond hair belong to the female group. Neural networks can capture this correlation, resulting in poor performance for certain subgroups (e.g., blond males) (Sagawa et al., 2019). We aim to ensure that the model does not capture spurious correlations in the data and accurately classifies samples from all subgroups. In particular, we measure the model’s bias using the worst-case group accuracy (WGA), i.e., group of samples where the spurious correlation does not apply. Several methods have been proposed to mitigate these biases in individual models (Kenfack et al., 2024a; Ye et al., 2024). When the group information is known, Sagawa et al. (2019) propose Group Distributionally Robust Optimization (DRO) that minimizes the loss of the group experiencing the maximum loss. However, group information can be costly to collect or unavailable due to privacy restrictions (Kenfack et al., 2024b). Several methods have been proposed in this setting to improve the worst-case group performance without group labels (Ye et al., 2024). For instance, Liu et al. (2021) and Nam et al. (2020) rely on a reference classifier to target and upweight worst-performing subgroups. These methods use the mistakes of the reference classifier to improve group robustness by up-weighting misclassified samples (Liu et al., 2021). The reference classifier is generally trained to amplify the misclassification of samples from the unknown worst-performing group (Nam et al., 2020). While these methods do not use group labels during the training, they require access to a small validation set with group labels for model selection and intensive hyperparameter tuning (Kenfack et al., 2024a).  Kirichenko et al. (2022) proposed Deep Feature Reweighting (DRF) for training group robust model by retraining the last layer of ERM model with a small subset of a held-out group-balanced dataset, which achieves state-of-the-art worst-case group performance. However, all debiasing methods use neural networks of high capacity, and they are less effective on models with smaller capacity Izmailov et al. (2022). Furthermore, the effect of distilling these debiased models on the student model’s robustness remains underexplored. This work builds on Kirichenko et al. (2022) for debiasing the teacher models by retraining their last layers using a held-out group balanced with fewer group annotations, and we investigate whether distilling knowledge from an ensemble of (debiased) teachers can lead to more robust student models.
@@
Several works have studied bias in knowledge distillation with a single teacher model (Lukasik et al., 2023; Lee & Lee, 2023; Lukasik et al., 2023; Tiwari et al., 2024; Bassi et al., 2024). In particular, Lukasik et al. (2023) demonstrates that teacher errors can be amplified by the student during distillation and proposed a mitigation strategy that distills only the confident predictions of the teacher. However, their study focuses on worst-class errors and KD with a single teacher, while we study worst-subgroup errors with multiple teachers. Lee & Lee (2023) propose an adapted version of Simple knowledge distillation (SimKD (Chen et al., 2022)) that transplants the last layer of the teacher to student and only distills features. With the teacher trained with Group DRO, they show that transplanting the teacher’s last layer to the student only improves the worst-case group performance if the feature distillation is performed by upweighting misclassified sample from a reference classifier (Lee & Lee, 2023). However, their method uses the reference classifier by Liu et al. (2021), where the efficiency of sample weighting requires intensive tuning of the number of epochs to train the reference classifier. Similarly,  Tiwari et al. (2024) uses the earlier layers of the neural network to train a reference classifier and shows that it improves the recall of worst group samples within the misclassification set, which are upweighted in the KD loss. However, their method also requires class labels to derive the misclassified samples. In Lukasik et al. (2023), the authors study where it is best to apply the debiasing mechanism (Group DRO) and conclude that applying the robust loss to both the teacher and student model improves the average performance along with the worst-case group performance. In contrast to these prior works, we study bias in knowledge distillation with multiple teachers without group information and class labels. We investigate whether the subgroup’s performance gain observed in deep ensemble models also applies when the knowledge of the ensemble is distilled to a single model. We aim to achieve better worst-case performance across subgroups when aggregating the outputs of multiple teachers in knowledge distillation. To the best of our knowledge, this represents the first study on bias in ensemble knowledge distillation.